# HyperGLM: HyperGraph for Video Scene Graph Generation and Anticipation

### v1

@@ Abstract
Multimodal LLMs have advanced vision-language tasks but still struggle with understanding video scenes. To bridge this gap, Video Scene Graph Generation (VidSGG) has emerged to capture multi-object relationships across video frames. However, prior methods rely on pairwise connections, limiting their ability to handle complex multi-object interactions and reasoning. To this end, we propose Multimodal LLMs on a Scene HyperGraph (HyperGLM), promoting reasoning about multi-way interactions and higher-order relationships. Our approach uniquely integrates entity scene graphs, which capture spatial relationships between objects, with a procedural graph that models their causal transitions, forming a unified HyperGraph. Significantly, HyperGLM enables reasoning by injecting this unified HyperGraph into LLMs. Additionally, we introduce a new Video Scene Graph Reasoning (VSGR) dataset featuring 1.9M frames from third-person, egocentric, and drone views and supports five tasks: Scene Graph Generation, Scene Graph Anticipation, Video Question Answering, Video Captioning, and Relation Reasoning. Empirically, HyperGLM consistently outperforms state-of-the-art methods across five tasks, effectively modeling and reasoning complex relationships in diverse video scenes.
@@ 1 Introduction
In this paper, we propose the Multimodal LLMs on a Scene HyperGraph (HyperGLM) approach to promote reasoning about multi-way interactions and higher-order relationships through a unified HyperGraph and LLMs, as illustrated in Fig. 1. To achieve this goal, we uniquely incorporate entity scene graphs, which capture spatial relationships between objects, with a procedural graph that models their causal transitions across video frames, forming a unified HyperGraph, as shown in Fig. 2. Significantly, our HyperGLM approach allows reasoning by injecting this unified HyperGraph into LLMs. In addition, we introduce a novel Video Scene Graph Reasoning (VSGR) dataset, comprising 1.9 million video frames surpassing existing benchmark datasets [18, 52, 47, 34, 35] in scale and annotation depth. Specifically, our VSGR dataset includes videos from third-person, egocentric, and drone perspectives and supports five tasks: Scene Graph Generation, Scene Graph Anticipation, Video Question Answering, Video Captioning, and Relation Reasoning. Notably, our VSGR dataset introduces a new Relation Reasoning task, setting it apart from existing video scene graph datasets, as shown in the comparison in Table 1.
@@ 4.1 Video Scene (Hyper)Graphs
where $P(r_{n}\mid r_{t}^{ij},v^{i,j})=w(r_{t}^{ij},r_{n})\times v^{i,j}$ is the probability of transitioning from relationship $r_{t}^{ij}$ to $r_{n}$, looking at object features. This allows the model to anticipate future interactions based on established temporal patterns observed.
@@ 5 Video Scene Graph Reasoning Dataset
Limitations of Current Datasets. Existing benchmarks (see Table 1) primarily support SGG and SGA, limiting their applicability on reasoning tasks. They suffer from inadequate support for VQA, VC, and RR, shallow annotations that fail to capture intricate object interactions, insufficient modeling of temporal dynamics, and poor multimodal integration. Thus, our VSGR dataset supports SGG, SGA, VQA, VC, and RR, enabling the reasoning capabilities of LLMs.
@@ 6.2 Ablation Study
Random Walk Parameters. The number of walks $(N_{w}$) and walk length $(N_{l}$) are introduced in Alg. 1, directly impacting the capacity to capture high-order relationships by determining the number of hyperedges. Although more hyperedges increase relational diversity, they can also introduce redundancy beyond an optimal point. Specifically, a higher $N_{w}$ broadens the range of sampled relationships, while a moderate $N_{l}$ balances depth. Our experiments demonstrate that optimal performances are achieved with 60 hyperedges ($N_{w}=60$ and $N_{l}=7$), shown in Fig. 6. Further experiments on the SGG task are provided in Table 3, and experiments on these parameters are included in the Appendices.
@@ 6.3 Comparison with State-of-the-Arts
Table 4: Comparison (%) on the VSGR and PVSG datasets for the Scene Graph Anticipation (SGA) task at Recall (R) and mean Recall (mR).
@@ 7 Conclusion
In this paper, we have presented HyperGLM, a novel approach to advance VidSGG. By injecting the information from the scene HyperGraph into LLMs, our approach has promoted accurate, context-rich interpretations of complex scenes, effectively reasoning about multi-way interactions and higher-order relationships. Extensive experiments on benchmarks, including PVSG, Action Genome, and our newly collected VSGR dataset, have proved that our HyperGLM approach outperforms state-of-the-art methods, demonstrating robustness and versatility across five tasks and video scenes, marking a notable advance in scene hypergraph-based understanding with language models.
@@
Future Work could explore an adaptive hypergraph-based representation that dynamically adjusts to evolving scene contexts. In addition, general world knowledge could be integrated into the procedural graph of the hypergraph combined with language models to improve context-aware reasoning.

### v2

@@ Abstract
Multimodal LLMs have advanced vision-language tasks but still struggle with understanding video scenes. To bridge this gap, Video Scene Graph Generation (VidSGG) has emerged to capture multi-object relationships across video frames. However, prior methods rely on pairwise connections, limiting their ability to handle complex multi-object interactions and reasoning. To this end, we propose the Multimodal Large Language Models (LLMs) on a Scene HyperGraph (HyperGLM), promoting reasoning about multi-way interactions and higher-order relationships. Our approach uniquely integrates entity scene graphs, which capture spatial relationships between objects, with a procedural graph that models their causal transitions, forming a unified HyperGraph. Significantly, HyperGLM enables reasoning by injecting this unified HyperGraph into LLMs. Additionally, we introduce a new Video Scene Graph Reasoning (VSGR) dataset featuring 1.9M frames from third-person, egocentric, and drone views and support five tasks. Empirically, HyperGLM consistently outperforms state-of-the-art methods, effectively modeling and reasoning complex relationships in diverse scenes.
@@ 1 Introduction
In this paper, we propose the Multimodal LLMs on a Scene HyperGraph (HyperGLM) approach to promote reasoning about multi-way interactions and higher-order relationships through a unified HyperGraph and LLMs, as illustrated in Fig. 1. To achieve this goal, we uniquely incorporate entity scene graphs, which capture spatial relationships between objects, with a procedural graph that models their causal transitions across video frames, forming a unified HyperGraph, as shown in Fig. 2. Significantly, our HyperGLM approach allows reasoning by injecting this unified HyperGraph into LLMs. In addition, we introduce a novel Video Scene Graph Reasoning (VSGR) dataset, comprising 1.9 million video frames surpassing existing benchmark datasets [18, 52, 47, 34, 35] in scale and annotation depth. Specifically, our VSGR dataset includes videos from third-person, egocentric, and drone perspectives. It supports five tasks: Scene Graph Generation, Scene Graph Anticipation, Video Question Answering, Video Captioning, and Relation Reasoning. Notably, our VSGR dataset introduces a new Relation Reasoning task, setting it apart from existing video scene graph datasets, as shown in the comparison in Table 1.
@@ 2 Related Work
In this section, we review advances in scene graph generation and hypergraph applications in computer vision, then discuss existing limitations and the advantages of our approach.
@@ 4.1 Video Scene (Hyper)Graphs
where $P(r_{n}\mid r_{t}^{ij},v^{i,j})=w(r_{t}^{ij},r_{n})\times v^{i,j}$ is the probability of transitioning from relationship $r_{t}^{ij}$ to $r_{n}$, looking at object features. This allows the model to anticipate future interactions based on established temporal patterns.
@@ 5 Video Scene Graph Reasoning Dataset
Limitations of Current Datasets. Existing benchmarks (see Table 1) primarily support SGG and SGA, limiting their applicability to reasoning tasks. They suffer from inadequate support for VQA, VC, and RR, shallow annotations that fail to capture intricate object interactions, insufficient modeling of temporal dynamics, and poor multimodal integration. Thus, our VSGR dataset supports SGG, SGA, VQA, VC, and RR, enabling the reasoning capabilities of LLMs.
@@ 6.2 Ablation Study
Hypergraph Parameters. The number of walks $(N_{w}$) and walk length $(N_{l}$) are introduced in Alg. 1, directly impacting the capacity to capture high-order relationships by determining the number of hyperedges. Although more hyperedges increase relational diversity, they can also introduce redundancy beyond an optimal point. Specifically, a higher $N_{w}$ broadens the range of sampled relationships, while a moderate $N_{l}$ balances depth. Our experiments demonstrate that optimal performances are achieved with 60 hyperedges ($N_{w}=60$ and $N_{l}=7$), shown in Fig. 6. Further experiments on the SGG task are provided in Table 3, and experiments on these parameters are included in the Appendices.
@@ 6.3 Comparison with State-of-the-Arts
Table 4: Comparison (%) on the VSGR and Action Genome datasets for the Scene Graph Anticipation (SGA) task at Recall (R) and mean Recall (mR).
@@ 7 Conclusion
In this paper, we have introduced HyperGLM, a novel VidSGG method that integrates scene hypergraph information into LLMs for context-aware and precise scene interpretation. Our approach effectively models complex interactions and higher-order relationships. It outperforms leading methods on benchmarks, including PVSG, Action Genome, and our newly collected VSGR dataset across five tasks.
@@
Acknowledgment. This material is based upon work supported by the National Science Foundation under Award No. OIA-1946391. We also acknowledge the Arkansas High-Performance Computing Center for providing GPUs.